1. Pollution is expensive to clean up. The cost to clean up p% of pollution is modeled by:
C=12000/(400-4p ) Where C is the cost of clean up in millions of dollars and p is the percentage of pollution.
A. How much will it cost to remove 80% of the pollution? My answer, 150 million
B. How much will it cost to remove 95% of pollution? My answer ,600 million
C. How much does it cost to remove 99% of pollution? My answer, 3,000 million or 3 billion
D.For what value(s) is this equation undefined?

Answer by Theo(13342)

since 400 - 4p is the denominator, it will be undefined when 400 - 4p = 0.
that will occur when p = 100% because 400 - 4*100 = 400 - 400 = 0.

you cannot divide by 0, so the answer, when the denominator is equal to 0, is undefined.

your other answers look ok.